GlueX TOF Meeting, September 20, 2016

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, September 20, 2016
11:00 am EDT
JLab: CEBAF Center, Room L207


  1. Announcements
    1. PMT order due in October 27
  2. Review of minutes from the previous meeting
  3. TOF Analysis Status
  4. Run Plan
  5. Calibration Updates
  6. NIM paper



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Thomas Britton, Brad Cannon, Paul Eugenio, Mark Ito (chair), Simon Taylor, Beni Zihlmann


The PMT order for 30 tubes is due in October 27. There is no immediate plan to replace tubes; these are spares.

Review of minutes from the previous meeting

We looked at the minutes from September 6.

  • TDC time windows. We mentioned last time that these need to be tightened up. Beni will tune the parameters once we have beam triggers.
  • New PMT bases with amplification. Mark paid a visit to Vladimir Popov to to discuss the possibility of using his design for the TOF. He was optimistic that it could be done. One concern of ours is the AC-coupled output. He says this is an essential feature of the design, but he thought that it should not be much of a problem even at our rates and that there may be other ameliorations to get around it. We need to meet with him to discuss the idea further.

High Voltage Scan Preparations

Beni had been working with Brad. They are transferring Beni's software and Brad is trying to reproduce the voltage settings that Beni finds. There are differences of 20 V with a spread of 20 V (Brad - Beni). He has also noticed differences in the fit parameters from the fit to gain vs. voltage. They are working to understand the differences.

Beni reminded us that we owe the Run Coordinator a one-pager describing the program we want to conduct during the run.

Measuring Difference Rate Between Upstream and Downstream Plane

Sasha has been looking for situations where the upstream plane is hit, but there is no corresponding hit in the downstream. He wants to look at how much we can reduce rate by putting up some plastic around the beamline. He is having a low yield of hits where he gets a four-fold coincidence: both ends, ADC and TDC. These are the objects that come out of the hit factory. He is exploring how to overcome the lack of statistics.

Non-Linear Calibration for the CAEN TDCs

Beni has succeeded in measuring the correction curve for the CAEN TDCs using a self trigger on the TOF. He can calibrate all channels in a few hours. He has not seen perfect reproducibility of the constants; there can be variation of a bin or two.

There was a discussion of whether we want to download the corrections to the TDCs or whether to turn off the on-board correction and do all corrections offline, using the calibration database. We have decided that an on-board correction would be much preferable, giving equivalent, improved resolution without having to keep track of boards and channels as a function of time. Ed is exploring whether this can be done; he needs instructions from the factory.

The correction cannot be read back to check it. On the other hand, we can re-do the calibration with the corrections loaded and check that the resulting spectrum is linear.

Paul pointed out that we should make sure Hall B is aware of this issue.

NIM Paper

Paul is working on it.